OpenStudy (anonymous):

A factory can produce two products, x and y, with a profit approximated by P = 14x + 22y – 900. The production of y can exceed x by no more than 100 units. Moreover, production levels are limited by the formula x + 2y ≤ 1,400. What production levels yield maximum profit?

OpenStudy (anonymous):

\[ \begin{array}{rcl} x + 2y &\leq& 1400 \\ y &\leq& x + 100 \end{array} \]You want to find the intersection points of the feasible region. Then plug each point into your profit equation and the point which give the highest value is the solution.

OpenStudy (anonymous):

Oh yeah, you also have: \[ x \geq 0 \\ y \geq 0 \]as making up your feasible region.

OpenStudy (anonymous):

thanks for the help.